# TIPO: Text to Image with Text Presampling for Prompt Optimization

### v2

@@ Appendix F Topic Distribution Visualization
By employing multi-level topic analysis, we empirically assessed whether the resulting topics were sufficiently large and diverse. This iterative process mitigates potential challenges such as substantial topic overlap, which can diminish distinctiveness when extracting a large number of topics [59].
To address the potential overlap and further assess the diversity and meaningfulness of the topics, we performed a secondary clustering [66]. We grouped the initially extracted topics into five major clusters using k-means clustering and evaluated the clustering performance by calculating the inertia (Sum of Squared Distances) [27], shown in Table 7, 8, and 9. Since the topics have already been filtered for meaningful content, a higher inertia value indicates greater diversity among the clusters, reflecting a broader range of valid and meaningful topics across the dataset. This two-tiered approach allows for a more nuanced analysis of topic diversity and ensures the robustness of the topic modeling against meaningless word groupings.
Size Promptdb GPT4o-mini Promptist TIPO
@@
Size Promptdb GPT4o-mini Promptist TIPO
@@
Size PromptDB GPT4o-mini Promptist TIPO
@@
We attach simple visualization of topics in scenery prompt generation, with topic n=100, cluster k=5.
@@
Figure 19: This visualization represents a filtered subset of posts from the Danbooru2023 dataset, centered on the ’scenery’ tag. The network graph is an ego network (depth 1), meaning it includes only nodes directly connected to the ’scenery’ tag. To refine the data and focus on meaningful associations, uncommon tags with fewer than 10 occurrences were excluded. The analysis, conducted using Gephi, focuses on nodes with a degree greater than 600 to highlight critical components. Nodes are color-coded by modularity class by Fast Unfolding Algorithm [10], revealing clusters of closely associated tags. Node size reflects Eigenvector Centrality[11], emphasizing tags that are both highly connected and influential within their network.

### v3

@@ Appendix H Topic Distribution Visualization
We empirically assessed whether the resulting topics were sufficiently large and diverse by employing multi-level topic analysis. This iterative process mitigates potential challenges such as substantial topic overlap, which can diminish distinctiveness when extracting a large number of topics [62].
To address the potential overlap and further assess the diversity and meaningfulness of the topics, we performed a secondary clustering [70]. We grouped the initially extracted topics into five major clusters using k-means clustering. We evaluated the clustering performance by calculating the inertia (Sum of Squared Distances) [27], shown in Table 10, 11, and 12. Since the topics have already been filtered for meaningful content, a higher inertia value indicates greater diversity among the clusters, reflecting a broader range of valid and meaningful topics across the dataset. This two-tiered approach allows for a more nuanced analysis of topic diversity and ensures the robustness of the topic modeling against meaningless word groupings.
Size MagicPrompt GPT4o-mini Promptist TIPO
@@
Size MagicPrompt GPT4o-mini Promptist TIPO
@@
Size MagicPrompt GPT4o-mini Promptist TIPO
@@
We attach a simple visualization of topics in scenery prompt generation, with topic n=100, cluster k=5.
@@
Figure 32: This visualization represents a filtered subset of posts from the Danbooru2023 dataset, centered on the ’scenery’ tag. The network graph is an ego network (depth 1), which includes only nodes directly connected to the ’scenery’ tag. To refine the data and focus on meaningful associations, uncommon tags with fewer than 10 occurrences were excluded. The analysis, conducted using Gephi, focuses on nodes with a degree greater than 600 to highlight critical components. Nodes are color-coded by modularity class by Fast Unfolding Algorithm [10], revealing clusters of closely associated tags. Node size reflects Eigenvector Centrality[11], emphasizing highly connected and influential tags within their network.
Appendix I Discussion and Future Work
Despite the promising performance demonstrated by TIPO, several areas merit further investigation:
Personalization and User Adaptation.
TIPO currently does not directly incorporate personalization methods (e.g. LoRA [30]). Incorporating such modules could further align generated outputs to individual user preferences, enabling more tailored creative assistance. Investigating joint training strategies that integrate personalized fine-tuning represents a valuable future direction.
Generalization to Uncommon Compositions.
While TIPO performs effectively on typical and broadly represented concepts, it may still struggle with highly unconventional combinations (e.g., “avocado chair”). Future research should systematically evaluate TIPO’s performance on rare or unusual prompts (but considered creative prompts), potentially exploring auxiliary training or adaptation methods to ensure broader compositional robustness.
Capacity and Backbone Scalability.
Given TIPO model’s architecture [66, 65], its capacity for handling highly complex or lengthy prompts could be inherently limited. Future work could investigate more extensive or specialized architectures enhance compositional and semantic capacity, ideally without compromising real-time inference efficiency, are crucial for practical deployment.